OpenStudy (anonymous):

Need help with Algebra 2!! What are the amplitude, period, and midline of f(x) = 5 sin(x − π) + 3? A. amplitude: 5; period: π; midline: y = 3 B. amplitude: 3; period: pi/2; midline: y = 5 C. amplitude: 3; period: pi/4; midline: y = 5 D. amplitude: 5; period: 2π; midline: y = 3

OpenStudy (anonymous):

@SolomonZelman Think you could help me out with this one too?

OpenStudy (solomonzelman):

I haven't done these in a while. I am replying instead of messaging, so that I can get a notification and view this later too. (you are probably thinking I am typing the answer, but well... -:( )

OpenStudy (solomonzelman):

Sorry

OpenStudy (anonymous):

It's fine! lol Thanks anyways though! (:

OpenStudy (solomonzelman):

@ganeshie8

ganeshie8 (ganeshie8):

\[ f(x) = 5 \sin(x − π) + 3\]

ganeshie8 (ganeshie8):

just keep in mind that \(-1 \le \sin (x) \le 1\)

ganeshie8 (ganeshie8):

that gives : \(-5 \le 5\sin (x) \le 5\)

ganeshie8 (ganeshie8):

so this wave goes swings up by 5 units, down by 5 units forever..

ganeshie8 (ganeshie8):

the \(+3\), shifts the wave up by 3 units as well

ganeshie8 (ganeshie8):

so the midline is : y = 3

ganeshie8 (ganeshie8):

the amplitude is : 5

ganeshie8 (ganeshie8):